For the following reaction, 4.82 grams of methane (CH4) are mixed with excess carbon tetrachloride . The reaction yields 48.2 grams of dichloromethane (CH2Cl2) .

What is the theoretical yield of dichloromethane (CH2Cl2) ?
What is the percent yield for this reaction ?
1
Expert's answer
2017-10-13T15:07:07-0400
The equation of the reaction is:
CH 4 + CCl 4  2CH 2 Cl 2
That means that 2 moles of CH 2 Cl 2 are obtained from 1 mole of CH 4 .
Molar mass of CH 4 is 16 g/mole.
Molar mass of CH 2 Cl 2 is 85 g/mole.
The amount of CH 4 is 4.82 (g) / 16 (g/mole) = 0,30125 (mole)
The theoretical amount of CH 2 Cl 2 is 2×0.30125 (mole) = 0.6025 (mole)
The theoretical yield of CH 2 Cl 2 is 0.6025 (mole) × 85 (g/mole) = 51.2125 (g)
The percent yield for the reaction is 48.2 (g) / 51.2125 (g) ×100% = 94.12%
Answer:
The theoretical yield of CH 2 Cl 2 is 51.2125 grams.
The percent yield for the reaction is 94.12%.